A 92.2 kg weight-watcher wishes to climb a
mountain to work off the equivalent of a large
piece of chocolate cake rated at 990 (food)
Calories.
How high must the person climb? The
acceleration due to gravity is 9.8 m/s
2
and 1
food Calorie is 103
calories.
Answer in units of km.
1
Expert's answer
2020-04-23T12:44:20-0400
"E=mgh"


"990 (4.184) (10^3) = 92.2\times 9.8\times h"


"h=4584 m = 4.58 km"